AN ELEVATOR THAT WEIGHS 3000 N IS ACCELERATED UPWARD AT 1 M/S2. WHAT FORCE DOES THE CABLE EXERT TO GIVE IT THIS ACCELERATION ?


1
Expert's answer
2022-02-18T08:23:06-0500

Given:

"mg=3000\:\rm N"

"a=1\:\rm m/s^2"


The tension in a cable

"T=ma+mg=mg(1+a/g)\\
=3000*(1+1/9.8)=3306\:\rm N"